rizza's age is one year more than twice norrie"s ge. eight years ago, the sum of thier ages was equal to rizza's age ten years from now. how old are they now?
Answer by Mathtut(3670)

ok equations to work with. first R(rizzas age)= 2 N(norries age)+1
now eight years ago the sum of their ages so lets write that as (R-8)+ (N-8)
= to rizzas age 10 years from now written as R+10
so we have R=2N+1 and R-8+N-8=R+10 substituting R=2N+1 into the 2nd equation we have (2N+1)-8+N-8=(2N+1)+10 solving for N we arrive at N=26 now plugging this value into R=2N+1 we arrive at R=53
so Rizzas age was 53 and Norries was 26
checking for eight years ago we have Rizza at 45 and Norrie at 18 add those 2 numbers together and you get 63 which is 10 years older than Rizza's age at this time.